# Monotonicity of Recurrence in Random Walks

### v2

@@ 2. A Counterexample to Monotonicity
In this section, we show that recurrence, positive recurrence, and transience are not monotonic properties of random walks on $\mathbb{N}^{2}$ with respect to the partial order $\preceq$. To do so, we construct a positive recurrent random walk $X$ and a transient random walk $Y$ such that $Y\preceq X$, proving Theorem 1.3.
@@ 2.1. Recurrent Random Walk
$\begin{cases}\text{right}&\text{w.p. }\frac{9}{14}\\ \text{up}&\text{w.p. }\frac{5}{14},\end{cases}\qquad\begin{cases}\text{up}&\text{w.p. }\frac{1}{2}\\ \text{right}&\text{w.p. }\frac{5}{12}\\ \text{left}&\text{w.p. }\frac{1}{12},\end{cases}\qquad\begin{cases}\text{down}&\text{w.p. }\frac{89}{120}\\ \text{up}&\text{w.p. }\frac{5}{24}\\ \text{right}&\text{w.p. }\frac{1}{20},\end{cases}\qquad\begin{cases}\text{up}&\text{w.p. }\frac{3}{8}\\ \text{down}&\text{w.p. }\frac{3}{8}\\ \text{left}&\text{w.p. }\frac{5}{24}\\ \text{right}&\text{w.p. }\frac{1}{24},\end{cases}$
@@ 2.2. Transient Random Walk
$\begin{cases}\text{right}&\text{w.p. }\frac{9}{14}\\ \text{up}&\text{w.p. }\frac{5}{14},\end{cases}\qquad\begin{cases}\text{up}&\text{w.p. }\frac{49}{100}\\ \text{right}&\text{w.p. }\frac{41}{100}\\ \text{left}&\text{w.p. }\frac{1}{10},\end{cases}\qquad\begin{cases}\text{down}&\text{w.p. }\frac{19}{25}\\ \text{up}&\text{w.p. }\frac{5}{25}\\ \text{right}&\text{w.p. }\frac{1}{25},\end{cases}\qquad\begin{cases}\text{up}&\text{w.p. }\frac{1}{100}\\ \text{down}&\text{w.p. }\frac{74}{100}\\ \text{left}&\text{w.p. }\frac{21}{100}\\ \text{right}&\text{w.p. }\frac{4}{100},\end{cases}$
@@ 3. Monotonicity in Homogeneous Random Walks
We now consider conditions under which recurrence and transience are indeed monotonic properties with respect to the partial order $\preceq$ and prove Theorem 1.5.
We will use the following notation. Let $\leq$ denote the partial order on $\mathbb{Z}^{2}$ given by $(i_{1},j_{1})\leq(i_{2},j_{2})$ if $i_{1}\leq i_{2}$ and $j_{1}\leq j_{2}$. For two elements $x=(i_{1},j_{1})$ and $y=(i_{2},j_{2})$ in $\mathbb{Z}^{2}$, we let $x+y$ denote $(i_{1}+i_{2},j_{1}+j_{2})$. We view $\mathbb{N}^{2}$ as a subset of $\mathbb{Z}^{2}$
@@ 4.1. A counterexample to monotonicity
$\begin{cases}\text{up}&\text{w.p. }0.65\\ \text{right}&\text{w.p. }0.33\\ \text{left}&\text{w.p. }0.01\\ \text{down}&\text{w.p. }0.01,\end{cases}\qquad\qquad\begin{cases}\text{right}&\text{w.p. }0.52\\ \text{up}&\text{w.p. }0.47\\ \text{left}&\text{w.p. }0.01,\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }0.49\\ \text{left}&\text{w.p. }0.48\\ \text{right}&\text{w.p. }0.03,\end{cases}$
@@
$-k-x_{0}+y_{0}\leq\lim_{n\to\infty}\mathbb{E}[S_{n}-S_{0}].$
As $|S_{1}-S_{0}|\leq 2$ always, we have
@@
$\begin{cases}\text{up}&\text{w.p. }0.01\\ \text{right}&\text{w.p. }0.32\\ \text{left}&\text{w.p. }0.02\\ \text{down}&\text{w.p. }0.65,\end{cases}\qquad\qquad\begin{cases}\text{right}&\text{w.p. }0.51\\ \text{up}&\text{w.p. }0.46\\ \text{left}&\text{w.p. }0.03,\end{cases}\qquad\qquad\begin{cases}\text{down}&\text{w.p. }0.50\\ \text{left}&\text{w.p. }0.49\\ \text{right}&\text{w.p. }0.01,\end{cases}$
@@
The $-2$ probability is tight when on the upper boundary, and the $2$ probability is tight when on the lower boundary. As $\mathbb{E}[B_{k}]>0$, our Hoeffding argument from Section 2.2 implies this random walk is transient.
@@ 4.2. Monotonicity in Homogeneous Random Walks
$\begin{cases}\text{left}&\text{w.p. }\ell_{q}\\ \text{right}&\text{w.p. }r_{q}\\ \text{up}&\text{w.p. }u_{q}\\ \text{down}&\text{w.p. }d_{q},\end{cases}\qquad\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{x}\\ \text{right}&\text{w.p. }r_{x}\\ \text{up}&\text{w.p. }u_{x},\end{cases}\qquad\qquad\begin{cases}\text{left}&\text{w.p. }\ell_{u}\\ \text{right}&\text{w.p. }r_{u}\\ \text{down}&\text{w.p. }d_{u},\end{cases}$
@@
Our proof is structurally similar to that of Theorem 1.5, so for sake of brevity and clarity we will omit additional elaboration when the reasoning is essentially identical to before. It suffices to show that we can couple $X$ and $Y$ such that $|j_{t}^{(2)}-j_{t}^{(1)}|+\max\{i_{t}^{(2)}-i_{t}^{(1)},0\}\leq 2\lceil k/2\rceil$ for all $t$. We couple $X$ and $Y$ in the same manner as before, where inward shifting only ever changes a right movement in $Y$ to a left movement. We couple the upper corner and boundary redirection rules as described in Definition 4.2, which implies that if one of $X$ and $Y$ is on the upper boundary and the other is at the upper corner, then ignoring inward shifting, if they (both) attempt to move up and the one on the upper boundary redirects right, the one at the upper corner will redirect right as well. We similarly couple the origin and lower boundary redirection rules.
We use the same notation of $Y_{t}^{\prime}=(i_{t}^{(2^{\prime})},j_{t}^{(2^{\prime})})$ as before. It suffices to show that for fixed $t\geq 1$, if $|j_{t-1}^{(2)}-j_{t-1}^{(1)}|+\max\{i_{t-1}^{(2)}-i_{t-1}^{(1)},0\}\leq 2\lceil k/2\rceil$, then $|j_{t}^{(2^{\prime})}-j_{t}^{(1)}|+\max\{i_{t}^{(2^{\prime})}-i_{t}^{(1)},0\}\leq 2\lceil k/2\rceil$, as inward shifting weakly decreases the second term. If neither $X$ nor $Y$ redirect, the desired inequality trivially holds by induction. Now, suppose there is exactly one redirection; we divide into casework on who redirects and the direction that the random walks attempted to go in.
Case 1: $X$ redirects trying to go up. Then $Y$ goes up, decreasing the first term by 1 from time $t-1$ to $t$, so the inequality holds regardless of how $X$ redirects, as its movement can only change one term by at most 1.
@@
Case 6: $Y$ redirects trying to go left. This implies $i_{t}^{(2^{\prime})}-i_{t}^{(1)}\leq 0$, so we trivially bound $|j_{t}^{(2^{\prime})}-j_{t}^{(1)}|+\max\{i_{t}^{(2^{\prime})}-i_{t}^{(1)},0\}\leq k+0\leq 2\lceil k/2\rceil$.
@@
Case 3: $X$ and $Y$ redirect trying to go left. As $X$ and $Y$ are both on the $y$-axis, by 2-periodicity they are an even distance apart. We have $\max\{i_{t-1}^{(2)}-i_{t-1}^{(1)},0\}=0$. If $k$ is odd, or if $k$ is even and they are at most $k-2$ apart, then $|j_{t-1}^{(2)}-j_{t-1}^{(1)}|\leq 2\lceil k/2\rceil-2$ and the sum of the two terms can increase by at most 2, yielding the desired inequality. Otherwise, $X$ and $Y$ are on the two corners. If either moves vertically then the first term decreases by 1 so the inequality holds regardless of how the other redirects; otherwise, they both move right, in which case the inequality continues to hold. ∎
@@
Following the reasoning of the proof of Theorem 1.6 and using the notation of the proof of Theorem 4.3, it suffices to couple $X$ and $Y$ such that $|j_{t}^{(1)}-j_{t}^{(2)}|+\max\{i_{t}^{(1)}-i_{t}^{(2)},0\}\leq 2\lceil k/2\rceil$ for all $t$. The coupling used is the same as that in the proof of Theorem 4.3, where inward shifting is replaced by outward shifting as in the proof of Theorem 1.6. The inductive verification that the coupling satisfies the desired inequality is identical to that in the proof of Theorem 4.3, where in all the casework we swap the roles of $X$ and $Y$. ∎

### v4

@@ 2. A Counterexample to Monotonicity
In this section, we show that recurrence, positive recurrence, and transience are not monotonic properties of random walks on $\mathbb{N}^{2}$ with respect to the partial order $\preceq$. To do so, we construct a positive recurrent random walk $Y$ and a transient random walk $X$ such that $X\preceq Y$, proving Theorem 1.3. In fact, in our example, the inequalities defining $X\preceq Y$ will all strictly hold.
Remark 2.1.
In a 1972 Russian paper [16] with English translation [15], as well as earlier in his 1970 book [14], Malyshev characterized the recurrent homogeneous random walks on $\mathbb{N}^{2}$, where in this context homogeneous means the mean drift is the same for all $\{(i,j):i,j\geq 1\}$, the mean drift is the same on the positive $x$-axis $\{(i,0):i\geq 1\}$, and the mean drift is the same on the positive $y$-axis $\{(0,j):j\geq 1\}$. His result applies to both our examples, i.e., it can be used to verify that $X$ is positive recurrent and $Y$ is transient. He did not comment on the non-monotonicity of his recurrence characterization, nor did he provide an example demonstrating monotonicity fails. Our proof of Theorem 1.3 differs from that of Malyshev, with our examples designed to be amenable to much simpler arguments, and we found it instructive to include this proof as we use a similar proof technique in the case of random walks on slabs, which have more complicated boundaries, in Section 4.
@@ 2.1. Recurrent Random Walk
[…]
@@ 2.2. Transient Random Walk
[…]
@@ 3. Monotonicity in Homogeneous Random Walks
We now consider conditions under which recurrence and transience are indeed monotonic properties with respect to the partial order $\preceq$ and prove Theorems 1.7 and 1.8. For simplicity of presentation, we prove Theorems 1.7 and 1.8 under an inward-homogeneous condition, and then in Remark 3.2 explain the minor modification needed to extend the proof to the full generality of weak inward-homogeneity.
We will use the following notation. Let $\leq$ denote the partial order on $\mathbb{Z}^{2}$ given by $(i_{1},j_{1})\leq(i_{2},j_{2})$ if $i_{1}\leq i_{2}$ and $j_{1}\leq j_{2}$. For two elements $x=(i_{1},j_{1})$ and $y=(i_{2},j_{2})$ in $\mathbb{Z}^{2}$, we let $x+y$ denote $(i_{1}+i_{2},j_{1}+j_{2})$. We view $\mathbb{N}^{2}$ as a subset of $\mathbb{Z}^{2}$.
@@
Remark 3.2.
Now assume $X$ is only weakly inward-homogeneous. Then when $X$ is on the positive $x$-axis, we can think of it as being lazy with probability $p_{x}:=\frac{r_{x}-r_{q}}{r_{x}}=1-\frac{r_{q}}{r_{x}}$, go right with probability $\frac{r_{q}}{r_{x}}r_{x}=r_{q}$, go left with probability $\frac{r_{q}}{r_{x}}\ell_{x}\geq\ell_{q}$, and go up with probability $\frac{r_{q}}{r_{x}}u_{x}\geq u_{q}$, so that its movement probabilities once it moves is still $\ell_{x}$, $u_{x}$, and $r_{x}$, as required. Similarly let $X$ be lazy on the positive $y$-axis. This laziness does not change whether $X$ is (positive) recurrent or transient. Then couple $(X,Y)$ as before with our three-stage process of picking one of the four directions, redirecting, and inward shifting, except now if say we need to redirect from choosing to go down while on the positive $x$-axis, with probability $\frac{p_{x}}{d_{q}}$ we do not move, and otherwise we redirect as necessary; of course, if $X$ and $Y$ are both on the positive $x$-axis, the decision to stay put or redirect is shared between $X$ and $Y$. Then it’s easy to verify that the inequalities $Y_{t}\leq X_{t}+s_{t}$ and $Y_{t}\geq X_{t}-s_{t}$ in the proofs of Theorems 1.7 and 1.8, respectively, still hold with this coupling, and so the proof follows identically.
@@ 4.1. A counterexample to monotonicity
[…]
@@
$k-x_{0}+y_{0}\leq\lim_{n\to\infty}\mathbb{E}[S_{n}-S_{0}].$
As $\left\lvert S_{1}-S_{0}\right\rvert\leq 2$ always, we have
@@
[…]
@@
The 2-2 probability is tight when on the upper boundary, and the $2$ probability is tight when on the lower boundary. As $\mathbb{E}[B_{k}]>0$, our Hoeffding argument from Section 2.2 implies this random walk is transient.
@@ 4.2. Monotonicity in Homogeneous Random Walks
[…]
@@
Our proof is structurally similar to that of Theorem 1.7, so for sake of brevity and clarity we will omit additional elaboration when the reasoning is essentially identical to before. It suffices to show that we can couple $X$ and $Y$ such that $\left\lvert j_{t}^{(2)}-j_{t}^{(1)}\right\rvert+\max\{i_{t}^{(2)}-i_{t}^{(1)},0\}\leq 2\lceil k/2\rceil$ for all $t$. We couple $X$ and $Y$ in the same manner as before, where inward shifting only ever changes a right movement in $Y$ to a left movement. We couple the upper corner and boundary redirection rules as described in Definition 4.2, which implies that if one of $X$ and $Y$ is on the upper boundary and the other is at the upper corner, then ignoring inward shifting, if they (both) attempt to move up and the one on the upper boundary redirects right, the one at the upper corner will redirect right as well. We similarly couple the origin and lower boundary redirection rules.
We use the same notation of $Y_{t}^{\prime}=(i_{t}^{(2^{\prime})},j_{t}^{(2^{\prime})})$ as before. It suffices to show that for fixed $t\geq 1$, if $\left\lvert j_{t-1}^{(2)}-j_{t-1}^{(1)}\right\rvert+\max\{i_{t-1}^{(2)}-i_{t-1}^{(1)},0\}\leq 2\lceil k/2\rceil$, then $\left\lvert j_{t}^{(2^{\prime})}-j_{t}^{(1)}\right\rvert+\max\{i_{t}^{(2^{\prime})}-i_{t}^{(1)},0\}\leq 2\lceil k/2\rceil$, as inward shifting weakly decreases the second term. If neither $X$ nor $Y$ redirect, the desired inequality trivially holds by induction. Now, suppose there is exactly one redirection; we divide into casework on who redirects and the direction that the random walks attempted to go in.
Case 1: $X$ redirects trying to go up. Then $Y$ goes up, decreasing the first term by 1 from time t1t-1 to $t$, so the inequality holds regardless of how $X$ redirects, as its movement can only change one term by at most 1.
@@
Case 6: $Y$ redirects trying to go left. This implies $i_{t}^{(2^{\prime})}-i_{t}^{(1)}\leq 0$, so we trivially bound $\left\lvert j_{t}^{(2^{\prime})}-j_{t}^{(1)}\right\rvert+\max\{i_{t}^{(2^{\prime})}-i_{t}^{(1)},0\}\leq k+0\leq 2\lceil k/2\rceil$.
@@
Case 3: $X$ and $Y$ redirect trying to go left. As $X$ and $Y$ are both on the $y$-axis, by 2-periodicity they are an even distance apart. We have $\max\{i_{t-1}^{(2)}-i_{t-1}^{(1)},0\}=0$. If $k$ is odd, or if $k$ is even and they are at most k2k-2 apart, then $\left\lvert j_{t-1}^{(2)}-j_{t-1}^{(1)}\right\rvert\leq 2\lceil k/2\rceil-2$ and the sum of the two terms can increase by at most 2, yielding the desired inequality. Otherwise, $X$ and $Y$ are on the two corners. If either moves vertically then the first term decreases by 1 so the inequality holds regardless of how the other redirects; otherwise, they both move right, in which case the inequality continues to hold. ∎
@@
Following the reasoning of the proof of Theorem 1.8 and using the notation of the proof of Theorem 4.3, it suffices to couple $X$ and $Y$ such that $\left\lvert j_{t}^{(1)}-j_{t}^{(2)}\right\rvert+\max\{i_{t}^{(1)}-i_{t}^{(2)},0\}\leq 2\lceil k/2\rceil$ for all $t$. The coupling used is the same as that in the proof of Theorem 4.3, where inward shifting is replaced by outward shifting as in the proof of Theorem 1.8. The inductive verification that the coupling satisfies the desired inequality is identical to that in the proof of Theorem 4.3, where in all the casework we swap the roles of $X$ and $Y$. ∎